OpenStudy (anonymous):

Lagrange Multipliers/Systems of Equations: find the points on the ellipse x^2 + 2y^2=1 where f(x,y)= xy has its extreme values........Ive gotten three equations in terms of lambda, y , and x, but I cant solve the System, can anyone provide me with a walkthrough?

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

this problem has my name in it, but can i solve it. I am very sleepy

OpenStudy (anonymous):

lets do this, i will help you solve for one variable, but then you well have to solve the rest, it will be easy because then you will only have to equations with only two unknown variables

OpenStudy (anonymous):

or maybe i will solve the whole thing

OpenStudy (anonymous):

what do you think?

OpenStudy (anonymous):

r u there? or did a robot ask this question?

OpenStudy (anonymous):

come on bud i dont want to solve this by myself

OpenStudy (anonymous):

srry

OpenStudy (anonymous):

was away for food

OpenStudy (anonymous):

man its 1:30 am in the morning i am so sleepy

OpenStudy (anonymous):

ya same here, i get...z=lambda y=z2x x=z4y

OpenStudy (anonymous):

and constraint x^2 + y^2 = 1

OpenStudy (anonymous):

but cant get to the books answer

OpenStudy (anonymous):

whats books answer? i just wanna make sure i am doing the problem right

OpenStudy (anonymous):

\[(\pm (1/\sqrt{2}, 1/2), (\pm (1/\sqrt{2}, -1/2)\]

OpenStudy (anonymous):

plus or minus only belongs to the x coordinate sorry

OpenStudy (anonymous):

(±1/2√,1/2),(±1/2√,−1/2)

OpenStudy (anonymous):

hmm, i get: